# Streamlined Supergravity

### v2

@@ 2 Construction of streamlined supergravity with a nilpotent superfield
We will sometimes call it “new” instead of “streamlined” for simplicity. The resulting bosonic action of scalars is
@@
The full supergravity action, including the fermions, was presented in the unitary gauge for local supersymmetry $\psi^{s}=0$ in Kallosh:2015sea , and a locally supersymmetric theory prior to gauge-fixing with a most general superpotential was given in Kallosh:2015tea . In the Appendix, we will give a short summary of the complete streamlined supergravity with scalars, fermions, and vectors following Kallosh:2015tea .
@@ 5 Discussion
In the Appendix, we present a complete streamlined supergravity including scalars, vectors, fermions, and one nilpotent chiral multiplet with a general superpotential $W(z^{i},s)=g(z^{i})+s\,f(z_{i})$ as constructed in Kallosh:2015tea . The vector part of the theory and a holomorphic gauge kinetic function $f_{AB}(z^{i},s)=f_{AB0}(z^{i})+s\,f_{ABs}(z^{i})$ and momentum maps ${\cal P}_{A}$ correspond to a standard supergravity, as given in the book Freedman:2012zz . We also have a comment there about the work in Cribiori:2017laj concerning Fayet-Iliopoulos terms in supergravity without gauged R-symmetry.
The major difference with the book Freedman:2012zz before gauge-fixing local supersymmetry is in the fermion sector of the theory, as one can see in eq. (43). In the unitary gauge where the fermion of the nilpotent multiplet $\psi^{s}$ vanishes, all extra terms vanish. The resulting theory is given in Eq. (45) for physical and nilpotent scalars, vectors, and fermions, where one has to put the nilpotent scalar to zero. This last step leaves the additional part in the scalar potential, $e^{K}F_{s}K^{s\bar{s}}\bar{F}_{\bar{s}}$, as compared with the standard supergravity without the nilpotent multiplet. This function is an arbitrary function of $(z^{i},\bar{z}^{\bar{i}})$.

### v3

@@ 2 Construction of streamlined supergravity with a nilpotent superfield
We will sometimes call it “new” rather than “streamlined” for simplicity. The resulting bosonic action of scalars is
@@
The full supergravity action, including the fermions, was presented in the unitary gauge for local supersymmetry $\psi^{s}=0$ in Kallosh:2015sea , and a locally supersymmetric theory prior to gauge-fixing with a most general superpotential was given in Kallosh:2015tea . In Appendix A, we will give a short summary of the complete streamlined supergravity with scalars, fermions, and vectors following Kallosh:2015tea . In Appendix B, we will describe the implementation of our methods to $SL(2,\mathbb{Z})$ invariant supergravity.
@@ 5 Discussion
In Appendix A, we present a complete streamlined supergravity including scalars, vectors, fermions, and one nilpotent chiral multiplet with a general superpotential $W(z^{i},s)=g(z^{i})+s\,f(z_{i})$ as constructed in Kallosh:2015tea . The vector part of the theory and a holomorphic gauge kinetic function $f_{AB}(z^{i},s)=f_{AB0}(z^{i})+s\,f_{ABs}(z^{i})$ and momentum maps ${\cal P}_{A}$ correspond to a standard supergravity, as given in the book Freedman:2012zz . We also have a comment there about the work in Cribiori:2017laj concerning Fayet-Iliopoulos terms in supergravity without gauged R-symmetry.
The major difference with the book Freedman:2012zz before gauge-fixing local supersymmetry is in the fermionic sector of the theory, as one can see in Eq. (43). In the unitary gauge where the fermion of the nilpotent multiplet $\psi^{s}$ vanishes, all extra terms vanish. The resulting theory is given in Eq. (45) for physical and nilpotent scalars, vectors, and fermions, where one has to put the nilpotent scalar to zero. This last step leaves the additional part in the scalar potential, $e^{K}F_{s}K^{s\bar{s}}\bar{F}_{\bar{s}}$, relative to the standard supergravity without the nilpotent multiplet. This function is an arbitrary function of $(z^{i},\bar{z}^{\bar{i}})$.
In Appendix B, we apply this general method to $SL(2,\mathbb{Z})$ invariant supergravity theories.
@@ Appendix A Complete streamlined supergravity with scalars, fermions, and vectors
Appendix B $SL(2,\mathbb{Z})$ invariant supergravity
We consider supergravity action with a physical chiral superfield where its scalar $\tau=iT$ is a modulus of the upper half-plane, and there is a nilpotent superfield with a scalar $s$ and fermion $\psi^{s}$. We study the action, including fermions, gravitino, and the fermion partner of $T$, which we call $\psi^{T}$, in the unitary gauge $\psi^{s}=0$. According to eq. (45), this means we have to start with standard supergravity with 2 chiral superfields with scalars $T,s$, and take into account that the potential (46) differs from the standard supergravity potential with only one superfield $T$.
Using the streamlined supergravity models with $W=W_{0}+s\,F_{s}$, one can construct a consistent supergravity embedding of the models with any potential $V(z^{i},\bar{z}^{\bar{i}})$. This method made it possible to construct a large set of supergravity models with $SL(2,\mathbb{Z})$- invariant potentials; see Sect. 9 of Kallosh:2024ymt . The Kähler invariant functional $G$ there is
$G(T,\bar{T};s,\bar{s})=-3\alpha\log(T+\bar{T})+\log|W_{0}+s\,F_{s}|^{2}+G_{s\bar{s}}(T,\bar{T})s\bar{s}\ ,$ (47)
where
$G_{s\bar{s}}=K_{s\bar{s}}={|F_{s}|^{2}\over(T+\bar{T})^{3\alpha}V(T,\bar{T})+3|W_{0}|^{2}(1-\alpha)}\ .$ (48)
The corresponding bosonic action is
Here, the scalar kinetic term is $SL(2,\mathbb{Z})$ invariant. Therefore, if the potential $V(T,\bar{T})$ is $SL(2,\mathbb{Z})$ invariant, the total bosonic part of the supergravity action is $SL(2,\mathbb{Z})$ invariant. It was sufficient here to use the simplest superpotential of the streamlined supergravity $W=W_{0}+s\,F_{s}$.
However, if we also want to have the fermion part of the action with gravitino and $\psi^{T}$ to be $SL(2,\mathbb{Z})$ invariant, we should use a more general $T$-dependent superpotential of the kind given in eq. (41). In supergravity with one chiral multiplet without the nilpotent one, the $SL(2,\mathbb{Z})$ invariant theory, including gravitino and $\psi^{T}$ was presented in Ferrara:1989bc . The proof of $SL(2,\mathbb{Z})$ invariance there required that the generating functional
$G(T,\bar{T})=-3\alpha\log(T+\bar{T})+\log W(T)+\log\bar{W}(\bar{T})$ (50)
should be invariant under the $SL(2,\mathbb{Z})$ transformations. The superpotential $W(T)$ must transform as a modular form of weight $3\alpha$ to compensate the $SL(2,\mathbb{Z})$ transformation of $K(T,\bar{T})=-3\alpha\log(T+\bar{T})$. The gravitino, which is neutral with respect to the $SL(2,\mathbb{Z})$ transformations, couples to $G(T,\bar{T})$.
$e^{G(T,\bar{T})\over 2}\bar{\psi}_{\mu R}\sigma^{\mu\nu}\psi_{\nu R}+h.c.$ (51)
The fermion superpartner of $T$, the field $\psi^{T}$, transforms as a modular form of weight 2-2 and it couples to $G(T,\bar{T})$ and its derivatives. This is why the $SL(2,\mathbb{Z})$ invariance of $G(T,\bar{T})$ is necessary and sufficient for the invariance of the fermion part of the action in the case studied in Ferrara:1989bc .
Now, as different from Ferrara:1989bc , we start the action with two multiplets: the physical one and the nilpotent one. Gravitino and $\psi^{T}$ are coupled to $G(T,\bar{T};s,\bar{s})$ and its $T$ derivatives, where $G(T,\bar{T};s,\bar{s})=G(T,\bar{T})+\dots$ Here $\dots$ means terms linear in $s$ and $\bar{s}$ and in $s\bar{s}$. But according to eq. (45) in the unitary gauge $\psi^{s}=0$, we have to take the standard action for both multiplets at $\psi^{s}=0,s=0$. Thus, in the final action, gravitino and $\psi^{T}$ are coupled to $G(T,\bar{T};s,\bar{s})\Big|_{s=0}=G(T,\bar{T})$ and its derivatives in $T,\bar{T}$ as shown in eq. (7) in Ferrara:1989bc .
To construct this $SL(2,\mathbb{Z})$ invariant supergravity with one physical multiplet and one nilpotent multiplet, we proceed as follows. We choose a superpotential
$W(T,s)=W(T)+s\,F_{s}\ ,$ (52)
where $W(T)$ is a modular form of weight $3\alpha$. After that, our generating potential (50) becomes $SL(2,\mathbb{Z})$ invariant. In this superpotential, the constant $W_{0}$ is replaced by a function $W(T)$, but we still keep the constant $F_{s}$. Our generating functional is
$G(T,\bar{T};s,\bar{s})=-3\alpha\log(T+\bar{T})+\log(W(T)+sF_{s})+\log(\bar{W}(\bar{T})+\bar{s}\bar{F}_{s})+K_{s\bar{s}}(T,\bar{T})s\bar{s}$ (53)
As we explained above, in the unitary gauge the physical fermions couple only to $G(T,\bar{T};s,\bar{s}))\Big|_{s=0}$, which is an expression given in eq. (50). Therefore, the total supergravity action, bosonic and fermionic part, is $SL(2,\mathbb{Z})$ invariant under the condition that the potential $V(T,\bar{T})$ is $SL(2,\mathbb{Z})$ invariant and eq. (46) is satisfied. We therefore find that for any choice of the $SL(2,\mathbb{Z})$ invariant potential $V(T,\bar{T})$ the inverse metric of the nilpotent superfield is
$K^{s\bar{s}}(T,\bar{T})={(T+\bar{T})^{3\alpha}V(T,\bar{T})-D_{T}W(T)K^{T\bar{T}}\bar{D}_{\bar{T}}\bar{W}(\bar{T})+3W(T)\bar{W}(\bar{T})\over F_{s}\bar{F}_{\bar{s}}}\ .$ (54)
The total Kähler potential is
$K(T,\bar{T},s,\bar{s})=-3\alpha\log(T+\bar{T})+{F_{s}\bar{F}_{\bar{s}}\over(T+\bar{T})^{3\alpha}V(T,\bar{T})-D_{T}W(T)K^{T,\bar{T}}\bar{D}_{\bar{T}}\bar{W}(\bar{T})+3W(T)\bar{W}(\bar{T})}s\,\bar{s}$
and the total superpotential is given in Eq. (52). These expressions for $K(T,\bar{T},s,\bar{s})$ and $W(T,s)$ define the supergravity theory, which is $SL(2,\mathbb{Z})$ invariant.